### v1

@@ A.4 Proof of Proposition 4
We now prove Proposition 4, which characterizes the equilibrium comovement between pass-through $\Phi_{ij}$ and the bilateral market shares $x_{ij}$ and $s_{ij}$.
Comparative Statics with Respect to $x_{ij}$
We begin by analyzing how pass-through $\Phi_{ij}$ responds to the importer’s buyer share $x_{ij}$. As discussed in the main text, $\Phi_{ij}$ depends on $x_{ij}$ through two channels: (i) the oligopsony markdown elasticity $\Gamma_{ij}^{\text{oligopsony}}$, which is non-monotonic (U-shaped) in $x_{ij}$, and (ii) the cost elasticity $\Lambda_{ij}$, which increases linearly in $x_{ij}$ when $\theta<1$.
When $\theta=1$, both elasticities are zero: $\Gamma_{ij}^{\text{oligopsony}}=\Lambda_{ij}=0$. Therefore, pass-through is constant in $x_{ij}$:
$\left.\frac{\partial\Phi_{ij}}{\partial x_{ij}}\right|_{\theta=1}=0.$
When $\theta<1$, however, both elasticities are active. The pass-through elasticity can be written as:
$\Phi_{ij}=\frac{1}{1+\Gamma_{ij}+\Lambda_{ij}}=\frac{1}{1+\cdots+\omega_{ij}\cdot\frac{\mu_{ij}^{\text{oligopsony}}}{\mu_{ij}}\cdot\Gamma_{ij}^{\text{oligopsony}}+\Lambda_{ij}},$
where $\cdots$” denotes terms that do not depend on $x_{ij}$. Differentiating with respect to $x_{ij}$:
$\frac{\partial\Phi_{ij}}{\partial x_{ij}}=-\frac{1}{(1+\Gamma_{ij}+\Lambda_{ij})^{2}}\cdot\frac{\partial}{\partial x_{ij}}\left(\omega_{ij}\cdot\frac{\mu_{ij}^{\text{oligopsony}}}{\mu_{ij}}\cdot\Gamma_{ij}^{\text{oligopsony}}+\Lambda_{ij}\right).$
Let:
$T_{1}\equiv\omega_{ij}\cdot\frac{\mu_{ij}^{\text{oligopsony}}}{\mu_{ij}}\cdot\Gamma_{ij}^{\text{oligopsony}},\quad T_{2}\equiv\Lambda_{ij}.$
We can expand these as:
@@
$T_{1}$ $=\omega_{ij}\cdot\left(\frac{(1-x_{ij})^{\frac{1}{\theta}-1}-\mu_{ij}^{\text{oligopsony}}}{\mu_{ij}}\right)(1-x_{ij})\varepsilon_{ij},$
$T_{2}$ $=\frac{1-\theta}{\theta}\cdot x_{ij}\cdot\varepsilon_{ij}.$
While $T_{1}$ is non-monotonic in $x_{ij}$, $T_{2}$ increases linearly in $x_{ij}$ and dominates $T_{1}$ for all $x_{ij}$. This can be verified numerically; for instance, when $\theta=\frac{1}{2}$:
$T_{1}+T_{2}=x_{ij}\cdot\varepsilon_{ij}\left(1-\frac{\omega_{ij}}{\mu_{ij}}\cdot\left(\frac{5}{2}-\frac{3}{2}x_{ij}\right)\right),$
which increases in $x_{ij}$ as the term in parentheses remains positive.
Hence, the term in the numerator of $\frac{\partial\Phi_{ij}}{\partial x_{ij}}$ increases in $x_{ij}$, implying:
$\operatorname{sign}\left(\frac{\partial\Phi_{ij}}{\partial x_{ij}}\right)=-\operatorname{sign}\left(\frac{\partial\Lambda_{ij}}{\partial x_{ij}}\right),$
so that:
$\frac{\partial\Phi_{ij}}{\partial x_{ij}}\leq 0.$
This proves that, under $\theta<1$, pass-through decreases in $x_{ij}$, a robust and testable implication of the model.
Comparative Statics with Respect to $s_{ij}$
We now turn to how pass-through $\Phi_{ij}$ varies with the exporter’s supplier share $s_{ij}$. Unlike the case of $x_{ij}$, this relationship is more complex, as all components, namely, $\Gamma_{ij}^{\text{oligopoly}}$, $\Gamma_{ij}^{\text{oligopsony}}$, $\Gamma_{ij}^{\omega}$, and $\Lambda_{ij}$, may vary with $s_{ij}$. Specifically,
$\Phi_{ij}=\frac{1}{1+\Gamma_{ij}+\Lambda_{ij}}=\frac{1}{1+\frac{(1-\omega_{ij})\mu_{ij}^{\text{oligopoly}}}{\mu_{ij}}\Gamma_{ij}^{\text{oligopoly}}+\frac{\omega_{ij}\mu_{ij}^{\text{oligopsony}}}{\mu_{ij}}\Gamma_{ij}^{\text{oligopsony}}+\left(1-\frac{\mu_{ij}^{\text{oligopoly}}}{\mu_{ij}}\right)\Gamma_{ij}^{\omega}+\Lambda_{ij}}.$
Differentiating with respect to $s_{ij}$ gives:
$\frac{\partial\Phi_{ij}}{\partial s_{ij}}$ $=-\frac{\partial}{\partial s_{ij}}\bigg{(}\frac{(1-\omega_{ij})\mu_{ij}^{\text{oligopoly}}}{\mu_{ij}}\Gamma_{ij}^{\text{oligopoly}}+\frac{\omega_{ij}\mu_{ij}^{\text{oligopsony}}}{\mu_{ij}}\Gamma_{ij}^{\text{oligopsony}}$
$\quad+\left(1-\frac{\mu_{ij}^{\text{oligopoly}}}{\mu_{ij}}\right)\Gamma_{ij}^{\omega}+\Lambda_{ij}\bigg{)}\cdot\left(1+\Gamma_{ij}+\Lambda_{ij}\right)^{-2}.$
When $\theta=1$, both $\Gamma_{ij}^{\text{oligopsony}}=0$ and $\Lambda_{ij}=0$, and the expression simplifies to:
$\frac{\partial\Phi_{ij}}{\partial s_{ij}}=-\frac{\partial}{\partial s_{ij}}\left(\frac{(1-\omega_{ij})\mu_{ij}^{\text{oligopoly}}}{\mu_{ij}}\Gamma_{ij}^{\text{oligopoly}}+\left(1-\frac{\mu_{ij}^{\text{oligopoly}}}{\mu_{ij}}\right)\Gamma_{ij}^{\omega}\right)\cdot(1+\Gamma_{ij})^{-2}.$
To build intuition, consider the limit $s_{ij}\to 1$, where $\omega_{ij}\to\phi$ and $\Gamma_{ij}^{\omega}\to 0$. Then:
@@
$\frac{\partial\Phi_{ij}}{\partial s_{ij}}\propto\begin{cases}-\frac{\partial}{\partial s_{ij}}\left(\frac{(1-\phi)\mu_{ij}^{\text{oligopoly}}}{\mu_{ij}}\Gamma_{ij}^{\text{oligopoly}}+\frac{\phi\mu_{ij}^{\text{oligopsony}}}{\mu_{ij}}\Gamma_{ij}^{\text{oligopsony}}+\Lambda_{ij}\right)&\text{if }\theta<1,\\ -\frac{\partial}{\partial s_{ij}}\left(\frac{(1-\phi)\mu_{ij}^{\text{oligopoly}}}{\mu_{ij}}\Gamma_{ij}^{\text{oligopoly}}\right)&\text{if }\theta=1.\end{cases}$
The term involving $\Gamma_{ij}^{\text{oligopoly}}$ is hump-shaped in $s_{ij}$, while the remaining components, particularly $\Lambda_{ij}$, decline monotonically with $s_{ij}$ due to $\frac{\partial\varepsilon_{ij}}{\partial s_{ij}}=-(\rho-\eta)<0$. Quantitatively, these latter terms dominate in most parameter ranges.
Thus, we have:
$\operatorname{sign}\left(\frac{\partial\Phi_{ij}}{\partial s_{ij}}\right)\propto\begin{cases}(+)&\text{if }\theta<1\text{ or }s_{ij}\text{ sufficientlylarge},\\ (-)&\text{if }\theta=1\text{ and }s_{ij}\text{ sufficiently small}.\end{cases}$
Due to the non-monotonicity and interaction of multiple channels, the relationship between $\Phi_{ij}$ and $s_{ij}$ does not yield a clean, general prediction. While pass-through may increase with $s_{ij}$ under decreasing returns to scale and sufficient buyer power, the sign of this relationship depends on the relative strength of strategic complementarities, oligopsony markdowns, and cost elasticity. As a result, we omit this relationship from Proposition 3 and focus instead on the more robust prediction involving $x_{ij}$.
Appendix B Theory: Discussion and Extensions
@@ B.1 Efficient Bargaining
In the efficient bargaining setup, the importer and exporter negotiate over a two-part tariff $(p,q)$ by maximizing a Nash product of generalized firm-specific gains from trade (GFT):
$\max_{p,q}\left[GFT_{ij}^{i}(p,q)\right]^{1-\phi}\left[GFT_{ij}^{j}(p,q)\right]^{\phi}\quad\text{s.t.}\quad GFT_{ij}^{i}\geq 0,\quad GFT_{ij}^{j}\geq 0$
where:
$GFT_{ij}^{i}(p_{ij},q_{ij})$ $\equiv\pi_{i}(p)-\tilde{\pi}_{i(-j)}=p_{ij}q_{ij}-\left(\theta c_{i}q_{i}-\theta\widetilde{c_{i}q_{i}}\right),$
$GFT_{ij}^{j}(p_{ij},q_{ij})$ $\equiv\pi_{j}(p)-\tilde{\pi}_{j(-i)}=\left(p_{j}q_{j}-\tilde{p}_{j}\tilde{q}_{j}\right)-p_{ij}q_{ij}.$
This formulation corresponds to a setting in which the firm pair first selects the input quantity $q_{ij}$ to maximize joint surplus, and then negotiates over the price $p_{ij}$ to determine how surplus is split.
@@
The efficient quantity solves:
$q_{ij}^{*}$ $:\quad\arg\max_{q}\left[GFT_{ij}^{i}(q)+GFT_{ij}^{j}(q)\right]$
$:\quad\frac{d(p_{j}q_{j})}{dq_{ij}}=\frac{d(\theta c_{i}q_{i})}{dq_{ij}}$
That is, the efficient quantity $q^{*}_{ij}$ equates the exporter’s marginal cost to the importer’s marginal revenue, similar to the outcome under vertical integration. Solving yields:
$q_{ij}^{*}:\quad p_{ij}=c_{i}(q_{ij}^{*}),$
so that the negotiated price equals the exporter’s marginal cost.
Given $q_{ij}^{*}$, the price $p_{ij}$ solves the Nash bargaining problem:
where $\phi\in(0,1)$ denotes the importer’s bargaining power. The terms $\Delta R_{j}(q_{ij})$ and $\Delta C_{i}(q_{ij})$ represent the additional revenue for the importer and the additional cost for the exporter attributable to the match:
$\Delta R_{j}(q_{ij})$ $=p_{j}q_{j}-\widetilde{p_{j}q_{j}},$
$\Delta C_{i}(q_{ij})$ $=\theta c_{i}q_{i}-\theta\widetilde{c_{i}q_{i}}.$
@@
$=(1-\phi)\cdot MC_{i}(q^{*}_{ij})+\phi\cdot AC(q^{*}_{ij})$
a weighted average of the per-unit downstream revenue gain and per-unit upstream cost increase from the match, with bargaining weights given by $\phi$.
Figure B.1 illustrates the set of feasible equilibria. The efficient quantity is $q^{*}$, while the negotiated price ranges between $MC(q^{*})$ and $AC(q^{*})$ depending on $\phi$.
@@ B.2 Supply-Driven Quantity Bargaining
as in equation (2.1). Solving the bargaining problem using derivations similar to those in Section A.1 of this Appendix yields the following expression for the bilateral price:
$p_{ij}=(1-\omega_{ij}^{S})\cdot c_{i}+\omega_{ij}^{S}\cdot\psi_{ij}^{-1}MRP_{j},$
@@
$\psi_{ij}=1+c^{\prime}_{i,q_{ij}}=1+\frac{1-\theta}{\theta}x_{ij}.$
Here, $\omega_{ij}^{S}$ is given by:
$\omega_{ij}^{S}\equiv\frac{\frac{\phi}{1-\phi}\cdot\lambda^{S}_{ij}}{1+\frac{\phi}{1-\phi}\cdot\lambda^{S}_{ij}}\quad\text{and}\quad\lambda^{S}_{ij}\equiv\frac{GFT_{ij}^{i}}{GFT_{ij}^{j}}\psi_{ij}=\frac{p_{ij}q_{ij}-\theta c_{i}q_{i}\Delta_{ij}^{x}}{p_{j}q_{j}\Delta_{ij}^{s}-p_{ij}q_{ij}}\psi_{ij}.$
@@
Hence, the bilateral price is a convex combination of the exporter’s marginal cost ($c_{i}$) and a term reflecting the markdown ($\psi_{ij}$) below the importer’s marginal revenue product ($MRP_{j}\equiv\frac{dp_{j}q_{j}}{dq_{ij}}$). As in the demand-driven quantity bargaining case, the weight $\omega_{ij}^{S}$ has an intuitive interpretation as the importer’s effective bargaining power.
For $\phi=0$, $\omega_{ij}^{S}=0$, and the allocation is efficient with $p_{ij}=c_{i}$. For $\phi=1$, $\omega_{ij}^{S}=1$, and we obtain the standard oligopsony benchmark where the input price is a markdown below the importer’s marginal revenue product, with the markdown equal to one plus the importer’s residual supply elasticity. This reflects the importer’s view, where the cost of increasing quantity is driven by raising the unit price for all inframarginal units when supply curves are upward-sloping.
Figure B.2 plots the set of feasible allocations. The red line indicates the pure oligopsony markdown, and the pink area represents the set of markdowns for $\phi\in(0,1)$.
While this case is intuitive, it is more complex to characterize due to the term $\lambda^{S}_{ij}$, which is itself a function of prices and quantities and thus no longer a simple function of market shares as in the baseline case. As a result, bringing this model to the data is less straightforward than our baseline model.
@@ B.3 Generalized Outside Option
In the baseline model, we assume that the importer’s (exporter’s) gains from trade are given by the firm’s total payoff from trading with all partners, minus the payoff from trading with all partners except exporter $i$ (importer $j$). In simpler terms, each importer (exporter) treats itself as the marginal buyer (supplier). This implies that, in case of disagreement, neither party considers the possibility of forming new relationships, which substantially influences the definition of outside profits for both importers and exporters.
We now consider a more general specification that imposes less structure on outside options. Let
$\Delta_{ij}^{c_{i}}=\frac{\tilde{c}_{i}}{c_{i}}\qquad\text{and}\qquad\Delta_{ij}^{c_{j}}=\frac{\tilde{c}_{j}}{c_{j}}$
@@
$GFT_{ij}^{i}(p_{ij})$ $=q_{ij}\left(p_{ij}-c_{i}\mu_{ij}^{\text{oligopsony},G}\right),$
@@
$\mu_{ij}^{\text{oligopsony},G}$ $:=\theta\left[\frac{\Delta_{ij}^{x}}{x_{ij}}\right],\qquad\Delta_{ij}^{x}:=\left[\frac{1-\Delta_{ij}^{c_{i}}(1-x_{ij})}{x_{ij}}\right],$
@@
$\mu_{ij}:=\frac{p_{ij}}{c_{i}}=(1-\omega^{G}_{ij})\mu_{ij}^{\text{oligopoly}}+\omega^{G}_{ij}\mu_{ij}^{\text{oligopsony},G},$ (B.1)
@@
$\omega^{G}_{ij}$ $:=\frac{\frac{\phi}{1-\phi}\lambda_{ij}^{G}}{1+\frac{\phi}{1-\phi}\lambda_{ij}^{G}}\in(0,1),$
$\lambda_{ij}^{G}$ $:=\frac{\gamma s_{ij}}{\Delta_{ij}^{s,G}}\cdot\frac{\nu-1}{\varepsilon_{ij}-1}>0.$
Equation (B.1) shares the structure of the markup equation in equation (2.6), but with two key differences. First, the term $\lambda_{ij}$ now depends on $\Delta_{ij}^{s,G}$, which cannot be expressed solely as a function of the supplier share. Second, the oligopsony markdown depends on $\Delta_{ij}^{x}$, which likewise cannot be written as a function of the buyer share alone.
To summarize, generalizing the structure of outside options preserves the overall structure of the markup equation and yields similar comparative statics. However, it introduces an identification problem: the effective bargaining power and oligopsony markdown now depend on unobserved terms, such as $\Delta_{ij}^{s,G}$ and $\Delta_{ij}^{x}$, that cannot be expressed solely as functions of market shares. In the absence of external data to estimate these terms, empirical implementation becomes infeasible without further assumptions. Imposing the more restrictive baseline assumption enables us to express both components as functions of observable bilateral market shares and a small set of structural parameters, allowing us to implement the model using available data and avoid excessive computational complexity.

### v2

@@ A.4 Proof of Proposition 4
Proposition 4 states that if the tariff pass-through elasticity $\Phi_{ij}$ decreases with the importer’s buyer share $x_{ij}$, then the exporter’s technology exhibits decreasing returns to scale, i.e. $\theta<1$.
Proof.
The bilateral pass-through elasticity is
$\Phi_{ij}=\frac{1}{1+\Gamma_{ij}+\Lambda_{ij}},$
where $\Gamma_{ij}$ and $\Lambda_{ij}$ denote the markup and marginal cost elasticities, respectively. Since $\partial\Phi_{ij}/\partial x_{ij}$ has the opposite sign of $\partial(\Gamma_{ij}+\Lambda_{ij})/\partial x_{ij}$, it suffices to show that
$\partial(\Gamma_{ij}+\Lambda_{ij})/\partial x_{ij}>0\quad\text{for}\quad\theta<1$
to prove the claim.
We first consider the case $\theta=1$.
Under constant returns to scale, the exporter’s marginal cost is constant and $\Lambda_{ij}=0$. Moreover, the oligopsony markdown equals one, implying $\Gamma^{\text{oligopsony}}_{ij}=0$. As $\Gamma_{ij}$ depends on $x_{ij}$ only through $\Gamma^{\text{oligopsony}}_{ij}$, it follows that $\partial\Gamma_{ij}/\partial x_{ij}=0$ and hence $\partial\Phi_{ij}/\partial x_{ij}=0$.
We now consider $\theta<1$.
The cost elasticity is
$\Lambda_{ij}=\frac{1-\theta}{\theta}x_{ij}\varepsilon_{ij},$
which is strictly increasing in $x_{ij}$ and independent of the bargaining parameter $\phi$. The markup elasticity depends on $x_{ij}$ only through the oligopsony component $\Gamma^{\text{oligopsony}}_{ij}$. Varying $\phi$ therefore only scales the contribution of this term. Consequently, if
$\left.\frac{\partial}{\partial x_{ij}}(\Gamma_{ij}+\Lambda_{ij})\right|_{\phi=1}>0,$
where the contribution of $x$ is maximized, then the same must hold for all $\phi\in[0,1]$. It is therefore sufficient to study the case $\phi=1$.
When $\phi=1$, the oligopsony markdown elasticity can be written as
@@
$\Gamma^{\text{oligopsony}}_{ij}=f(x_{ij})\,\varepsilon_{ij},\qquad f(x)=\left(\frac{x(1-x)^{\frac{1}{\theta}-1}}{\theta\left[1-(1-x)^{\frac{1}{\theta}}\right]}-1\right)(1-x).$
Hence,
@@
$\frac{\partial}{\partial x_{ij}}\bigl(\Gamma^{\text{oligopsony}}_{ij}+\Lambda_{ij}\bigr)=\varepsilon_{ij}\left[f^{\prime}(x_{ij})+\frac{1-\theta}{\theta}\right].$
The function $f$ is smooth on $(0,1)$ and U-shaped: $f^{\prime}(x)<0$ for small $x$, $f^{\prime}(x)>0$ for large $x$, and $f^{\prime}(x)$ is strictly increasing in $x$. Thus the minimum of $f^{\prime}(x)$ over $(0,1)$ is attained at the boundary as $x\to 0$.
A first-order expansion around $x=0$ yields
$f(x)=\frac{\theta-1}{2\theta}x+O(x^{2}),\qquad\lim_{x\to 0}f^{\prime}(x)=\frac{\theta-1}{2\theta}.$
When $\theta<1$, this implies
$f^{\prime}(x)\geq-\frac{1-\theta}{2\theta}>-\frac{1-\theta}{\theta}\qquad\forall x\in(0,1).$
Therefore,
$f^{\prime}(x_{ij})+\frac{1-\theta}{\theta}>0\quad\forall x_{ij}\in(0,1),$
and hence
$\frac{\partial}{\partial x_{ij}}\bigl(\Gamma_{ij}+\Lambda_{ij}\bigr)>0\qquad\forall\phi\in[0,1]\quad\&\quad\theta<1.$
Appendix B Theory: Extensions
@@ B.1 Efficient Bargaining
In the efficient bargaining setup, the importer and exporter jointly choose the price $p_{ij}$ and quantity $q_{ij}$ by maximizing the Nash product:
$\max_{p_{ij},\,q_{ij}}\left(GFT_{ij}^{i}\right)^{1-\phi}\left(GFT_{ij}^{j}\right)^{\phi}\quad\text{s.t.}\quad GFT_{ij}^{i}\geq 0,\quad GFT_{ij}^{j}\geq 0,$
where $GFT_{ij}^{k}\equiv\pi^{k}_{ij}-\tilde{\pi}^{k}_{(-\cdot)}$, $k\in\{i,j\}$, are as defined in Section 2. This corresponds to a setting in which the firm pair first selects $q_{ij}$ to maximize joint surplus, and then negotiates over $p_{ij}$ to split it.
The first-best quantity is determined where the exporter’s marginal cost equals the importer’s marginal revenue product
$q_{ij}^{*}:MR_{j}(q^{*}_{ij})=MC_{i}(q^{*}_{ij}).$
Given $q_{ij}^{*}$, the price $p_{ij}$ solves:
where $\Delta R_{j}(q_{ij})\equiv R_{j}(q_{j})-\tilde{R}_{j}(\tilde{q}_{j})$ and $\Delta C_{i}(q_{ij})\equiv TC_{i}(q_{i})-TC_{i}(\tilde{q}_{i})$ are the incremental revenue for the importer and the incremental cost for the exporter from the match.
@@
a weighted average of the per-unit downstream revenue gain and per-unit upstream cost increase from the match, with bargaining weights given by $\phi$. Figure B.1 illustrates this result. The efficient quantity $q^{*}$ is determined where the exporter’s marginal cost $\mathrm{MC}(q_{i})$ equals the importer’s marginal revenue product $p(q_{i})$. At $q^{*}$, the negotiated price lies between $\Delta C_{i}/q^{*}=\mathrm{AC}(q^{*})$ and $\Delta R_{j}/q^{*}=p(q^{*})$. The weight $\phi$ determines the exact price within this interval: as $\phi\to 0$ the price approaches $\Delta R_{j}/q^{*}$, and as $\phi\to 1$ it approaches $\Delta C_{i}/q^{*}$.
@@ B.2 Supply-Driven Quantity Bargaining
as in equation (2.1). Since $p_{ij}=MC_{i}(q_{ij})$ holds on the supply curve, solving the Nash bargaining problem yields the following first-order condition characterizing the equilibrium quantity (and hence price):
$\psi_{ij}\,p_{ij}-MRP_{j}\;=\;\frac{1-\omega_{ij}^{S}}{\omega_{ij}^{S}}\,(\psi_{ij}-1)\,p_{ij},$
@@
$\psi_{ij}=1+c^{\prime}_{i,q_{ij}}=1+\frac{1-\theta}{\theta}x_{ij},$
and $\omega_{ij}^{S}$ has the same functional form as $\omega_{ij}$ in Proposition 1:
$\omega_{ij}^{S}\equiv\frac{\phi\,\psi_{ij}\,GFT_{ij}^{i}}{(1-\phi)\,GFT_{ij}^{j}+\phi\,\psi_{ij}\,GFT_{ij}^{i}}\in(0,1),$
where $GFT_{ij}^{i}=p_{ij}q_{ij}-\theta MC_{i}q_{i}\Delta_{ij}^{x}$ and $GFT_{ij}^{j}=p_{j}q_{j}\Delta_{ij}^{s}-p_{ij}q_{ij}$ are as in Section 2.
The formula has the same structure as $\omega_{ij}$ in Proposition 1: $\omega^{S}_{ij}$ is a ratio of the importer’s weighted gains $\phi\,\psi_{ij}\,GFT^{i}_{ij}$ to total weighted gains, where the supply-slope factor $\psi_{ij}$ plays the role of $1+e_{i}$ (the labor-supply elasticity factor) in their model. The key difference from the baseline is that $\omega^{S}_{ij}$ depends on equilibrium gains-from-trade levels rather than on observable market shares alone, which is why the equilibrium condition does not reduce to a closed-form convex combination.
For $\phi=0$, $\omega_{ij}^{S}=0$: the exporter captures the full importer surplus ($GFT_{ij}^{j}=0$), and the equilibrium price is pinned by $\Delta R_{j}(q_{ij})=p_{ij}\,q_{ij}$. For $\phi=1$, $\omega_{ij}^{S}=1$: the first-order condition collapses to $\psi_{ij}\,p_{ij}=MRP_{j}$, recovering the standard oligopsony benchmark where the input price is a markdown factor $\psi_{ij}^{-1}$ below the importer’s marginal revenue product. Here $\psi_{ij}=1+$ the inverse of the importer’s residual supply elasticity, reflecting the cost of raising the unit price on all infra-marginal units when the supply curve is upward-sloping.
@@
Figure B.2 illustrates the equilibrium. The blue curve is the exporter’s marginal cost $MC_{i}$ (equivalently, the supply curve), and the black curve is the importer’s marginal revenue product $MRP_{j}$. The red curve is the importer’s marginal factor cost $\psi_{ij}\,MC_{i}$, which exceeds $MC_{i}$ whenever $\theta<1$. At $\phi=1$ the equilibrium lies at point A, where $MRP_{j}=\psi_{ij}\,MC_{i}$: the importer acts as a pure monopsonist, restricting quantity below the efficient level $q^{*}$ (where $MC_{i}=MRP_{j}$). As $\phi$ decreases, bargaining power shifts toward the exporter, moving the equilibrium up the supply curve toward $q^{*}$. The shaded region on the supply curve indicates the feasible price range for $\phi\in(0,1)$.
A key implication of this protocol is that prices and quantities positively co-move in equilibrium. Because the equilibrium always lies on the exporter’s supply curve $p_{ij}=MC_{i}(q_{ij})$, which is upward-sloping under decreasing returns ($\theta<1$), any shock that increases the equilibrium quantity also raises the equilibrium price. This stands in sharp contrast to the demand-driven baseline, where the equilibrium lies on the importer’s downward-sloping demand schedule and price increases are accompanied by quantity reductions. The data support the demand-driven prediction: tariff-induced price increases are accompanied by quantity declines ($p\uparrow,\,q\downarrow$).
@@ B.3 Generalized Outside Option
In the baseline model, the importer’s (exporter’s) gains from trade are the difference between its payoff when trading with all current partners and its payoff when trading with all partners except exporter $i$ (importer $j$). This rules out two forces that may matter in richer environments: after a failed negotiation, firms may form or reallocate relationships, and broader equilibrium adjustments may shift disagreement payoffs by changing costs or demand elsewhere in the network.
We now consider a more general specification that places less structure on outside options and allows these effects to enter through disagreement payoffs. For tractability, we assume $\varrho=1$ throughout this section (constant returns to scale for the importer), so that $\eta-1=(\nu-1)\gamma$. Let
$\Delta_{ij}^{MC_{i}}=\frac{\widetilde{MC}_{i}}{MC_{i}}\qquad\text{and}\qquad\Delta_{ij}^{c_{j}}=\frac{\widetilde{MC}_{j}}{MC_{j}}$
@@
$GFT_{ij}^{i}(p_{ij})$ $=q_{ij}\left(p_{ij}-MC_{i}\mu_{ij}^{\text{oligopsony},G}\right),$
@@
$\mu_{ij}^{\text{oligopsony},G}$ $:=\theta\left[\frac{\Delta_{ij}^{x}}{x_{ij}}\right]\qquad\text{with}\qquad\Delta_{ij}^{x}:=\left(1-\Delta_{ij}^{MC_{i}}(1-x_{ij})\right)$
and
@@
$\mu_{ij}:=\frac{p_{ij}}{MC_{i}}=(1-\omega^{G}_{ij})\mu_{ij}^{\text{oligopoly}}+\omega^{G}_{ij}\mu_{ij}^{\text{oligopsony},G},$ (B.1)
@@
$\omega^{G}_{ij}=\frac{\phi\,\xi^{G}(GFT_{ij}^{j},p_{ij})}{\phi\,\xi^{G}(GFT_{ij}^{j},p_{ij})+(1-\phi)(\varepsilon_{ij}-1)}\in(0,1).$
Here
$\xi^{G}(GFT_{ij}^{j},p_{ij})\equiv-\frac{d\ln GFT_{ij}^{j}}{d\ln p_{ij}}=\frac{(\eta-1)s_{ij}}{\Delta_{ij}^{s,G}}$
is the generalized elasticity of the importer’s gains from trade with respect to the bilateral price. This expression is equivalent to the baseline elasticity up to the denominator. In the baseline model, that denominator is pinned down by the supplier share alone. Here, instead, it also depends on how the importer’s disagreement payoff changes after a failed negotiation, as summarized by $\Delta_{ij}^{s,G}$. Since $\Delta_{ij}^{s,G}$ reflects unobserved changes in marginal cost and outside options, effective bargaining power is no longer identified from $s_{ij}$ alone.
Equation (B.1) shares the weighted-average structure of equation (2.10), but with two key differences. First, the generalized elasticity $\xi^{G}(GFT_{ij}^{j},p_{ij})$ now depends on $\Delta_{ij}^{s,G}$, so effective bargaining power is no longer pinned down by the supplier share alone. Second, the oligopsony markdown depends on $\Delta_{ij}^{x}$, so it is no longer determined by the buyer share alone.
The main implication is that outside options can absorb both extensive-margin reallocation and broader equilibrium forces, but once they do, the pricing equation depends on unobserved objects. In particular, $\Delta_{ij}^{s,G}$ and $\Delta_{ij}^{x}$ are not recoverable from bilateral market shares alone. Without additional data or stronger assumptions, the model is therefore not identified in a way that is useful for our empirical exercise. The baseline specification imposes more structure precisely to recover markup components from observable bilateral shares and a small set of parameters.